Julie is a small farm owner. She is endowed with 30 Kilos of Bananas and 10 Kilos of apples, which are her only source of income. She consumes bananas and apples in fixed proportions. One kilo of bananas per one kilo of apples. She can trade both fruits in a market without any transactions costs. The market price of each of fruits is $10. To maximize her utility, Julie will (select one)

1. not trade any fruit

2. sell 1 kilo of bananas and buy 1 kilo of apples

3. sell 10 kilos of bananas and buy 10 kilos of apples

4. sell 10 kilos of apples and buy 10 kilos bananas

Second part f the question: With the initial endowment, the market price of bananas increases to $30 per kilo, while the price of apples stays unchanged. After the price change, Julie will

1. Not Trade any fruit

2. Sell 5 kilos of Bananas and buy 15 kilos of apples

3. Sell 10 kilos of Bananas a buy 30 kilos of apples

4. Sell 15 Kilos of bananas and buy 45 kilos of apples
Answer #1

Let Bananas be B and apples be A.

1.

Julie's budget is given as:

10B + 10A = 10 x 30 + 10 x 10 = 400

With fixed proportion of 1:1, in equilibrium we have:

A = B

Using this in budget, we get:

A = B = 20

Ans: 3. sell 10 kilos of bananas and buy 10 kilos of apples.

2.

Julie's budget is given as:

30B + 10A = 30 x 30 + 10 x 10 = 1000

With fixed proportion of 1:1, in equilibrium we have:

A = B

Using this in budget, we get:

A = B = 25

Ans: 2. Sell 5 kilos of Bananas and buy 15 kilos of apples.

If you are satisfied with the answer, please provide a positive rating. Feel free to comment in case of queries.

Have a nice day ahead!